The position of a particle moving along X axis varies with time as X=10-2t
1
Expert's answer
2021-09-10T12:18:31-0400

The velocity of the particle is the first derivative by time, or


"v=\frac{dx}{dt}=\frac{d}{dt}(10-2t)=-2\text{ m/s}."